OpenStudy (anonymous):

If m = n + 3, then a^m/a^n

OpenStudy (amistre64):

a^(n+3) ------- = a^(n+3 - (m-3)) a^(m-3)

OpenStudy (amistre64):

n+3-m+3 = n-m +6 n-(n+3)+6 n-n-3+6 = 3 a^3

OpenStudy (anonymous):

ok and it has a 2nd part the only thing that changes is the number is 4 instead of 3

OpenStudy (amistre64):

m = n+4 right? im gonna guess and say a^4 :) but lets see if im right lol n+4-m+4 = n-m +8 n-(n+4)+8 n-n-4+8 = 4 a^4...yes!!!

OpenStudy (anonymous):

thank you you have been so helpful!!!

OpenStudy (amistre64):

:) youre welcome